# Funding AI for Good: A Call for Meaningful Engagement

### v2

@@ 5.2.1. Background: situating AI4SG in well-scoped contexts
Funding calls should orient towards projects that address well-scoped and contextually relevant problems. This doesn’t mean that funders should pre-define the problems, but they can support applicants by offering examples of relevant AI use cases or existing literature review on the social domain (Section 4.1.2). It may also involve requiring the active participation of communities during the problem formulation and solution brainstorming stage, where applicants and communities collaboratively investigate which technologies are appropriate for their specific problem. In a systematic review of CCA, Cooper et al. highlight that many projects involve communities only during limited stages (e.g., delivery of activities), rather than throughout the full research process—including problem framing and data analysis (Cooper et al., 2022). Yet, as Bratteteig and Wagner argue, engaging communities in deciding which problems to address is key, not only for inclusion but for redistributing power and ensuring they have real influence over project direction and outcomes (Bratteteig and Wagner, 2016). This level of participation reflects a deeper form of agency than simply evaluating or implementing pre-defined solutions (Delgado et al., 2023).
@@ 5.2.2. Expected Outcomes: defining success with intended communities
Successful outcomes in AI4SG projects should be defined collaboratively with the communities they aim to serve. Community perspectives on success often differ from funders’ or researchers’ goals, which tend to emphasize metrics like model accuracy or academic output (Lin et al., 2024; Erete et al., 2023)Lin et al. (2024) found that community organizations in AI4SG partnerships value learning about AI innovations, technical capacity building, and visibility of the projects even when real-world deployment isn’t achieved (Lin et al., 2024). Relatedly, Kumar and Dell (2018) report that development organizations prioritize accessible communication over formal research publications. Aligning success metrics early on helps manage expectations and reduces the risk of disappointment as projects progress (Taylor et al., 2013). While stakeholder goals may still sometimes conflict (Ho et al., 2009a), projects should aim for mutual benefit—for instance, through fair compensation for communities’ time and expertise or by supporting informal learning opportunities (Carroll and Rosson, 2013). This would also require attention to non-traditional (and often not quantifiable) metrics that capture relationship quality (Cooper et al., 2022; Erete et al., 2023). For example, funders could request letters of support and reports of experiences from communities as part of the project’s evaluation, alongside the quantitative evaluation of AI systems.
When real-world deployment is the goal, it is essential to provide funding support and maintenance resources post-deployment, which is notably absent from most funding programs in our dataset. In categorizing ICTD projects’ outcomes, Heeks calls attention to the high rate of projects that might have initially succeeded but break down over time (Heeks, 2002). As Jackson argues, “repair and maintenance” is a critical part of technology design, yet it pervasively falls out of view when the dominant discourse of technology focuses on innovations as a production, and not a process (Jackson, 2014). The maintenance work of a technology system is not only necessary for sustaining the system that will inevitably keep breaking (Jackson, 2014), it also builds relationships among workers who collaborate on the practical work (Irani and Silberman, 2014; D’Ignazio, 2024). Sustainable maintenance can involve building local capacity and intentionally transferring technology ownership to communities. For example, the research team behind SignSupport, a mobile app co-developed with the Deaf Community of Cape Town to support pharmacy interactions, trained community members to manage a computer lab and take on system administration roles, while researchers stepped back to allow community leadership to grow without interference (Dearden and Tucker, 2016).
@@
Funding calls should be structured to include intended community members as co-leaders or co-applicants, ensuring they have the power to shape project direction from the outset—and importantly, that they receive funding directly. As Bratteteig and Wagner point out, the “skill composition” of a design team significantly shapes how problems and solutions are defined (Bratteteig and Wagner, 2016). Including community members in core project roles enables them to define the problem space and success criteria, and ensures their labor and expertise are properly acknowledged and compensated (Pine et al., 2020). This approach also helps redistribute power within AI4SG collaborations, supporting more equitable relationships between communities and research institutions (Erete et al., 2023). Crucially, when engaging communities throughout the project, it is important to recognize and compensate for the various forms of labor they contribute—from sharing lived experience and domain expertise to facilitating connections within their communities. Future HCI research could work towards quantifying both visible and invisible labor (see (Toxtli et al., 2021) and (Li et al., 2022)), while also acknowledging that some forms of labor in community engagement, such as emotional labor, may be difficult or impossible to measure.
@@ 5.2.4. Funders’ Support: valuing community engagement as a necessary part of AI4SG work
Funders can better support community engagement in AI4SG by recognizing that relationship-building is a foundational phase of the work, which often occurs before a clear project direction is even established (Carroll and Rosson, 2013; Cooper et al., 2022). This phase is essential to community-collaborative approaches (Cooper et al., 2022), and it requires the routine presence of project teams, which takes time and emotional labor (Dearden and Tucker, 2016). Long-term relationships enable shared agendas to emerge, with community priorities guiding technology development from the outset (Dearden and Tucker, 2016; Carroll and Rosson, 2013). One concrete step funders can take is to extend the typical AI4SG funding timeline (in our corpus: with a median of one and a half years) to accommodate this early-stage work. Relationship-building alone can take months (Erete et al., 2023) or even years (Carroll and Rosson, 2013; Dearden and Tucker, 2016), and should be supported as a formal project phase, even prior to defining problems or solutions. One key challenge is determining appropriate timelines and support for this work. Future research in CCA could analyze researchers’ reflections (e.g., (Erete et al., 2023; Dearden and Tucker, 2016; Carroll and Rosson, 2013)) to better understand the duration and nature of relationship-building across projects.
@@
Furthermore, funders should also treat community engagement as a critical skill that requires training and support, on par with technical expertise. Many AI practitioners are trained in Computer Science departments that don’t routinely provide training in community engagement. As Okolo and Lin observe, AI4SG practitioners often use a hodgepodge of ad-hoc approaches to engage with end users (Okolo and Lin, 2024). Funders should recognize this knowledge gap and provide clear guidance to project teams on how to properly identify stakeholders and engage with end beneficiaries. There are multiple places where this guidance could be provided: as part of the funding call itself or additional training support that many funders already provide. Funders could consider citing existing guides in community engagement from the HCI community, such as the Equity-Centered Community Design Field Guide (Creative Reaction Lab, 2024), the Atlanta Community Engagement Playbook (Taranji Alvarado et al., 2024), and “readiness for partnership” checks (Pine et al., 2020). Other design toolkits like the Envisioning Cards (Friedman et al., 2024) might also be useful for AI4SG project teams to identify and engage with stakeholders. Since AI introduces additional hurdles to community-based designs, including requiring intensive data work (Miceli and Posada, 2022; Passi and Sengers, 2020; Sambasivan et al., 2021) and transparency issues (Okolo and Lin, 2024), HCI researchers could support funders by tailoring existing toolkits to the specific contexts of AI4SG, building on emerging Participatory AI4SG frameworks like the PACT framework (Bondi et al., 2021).
@@ 5.3. Calling for Meaningful Engagement: Recommendations for the Funding Design Process
Beyond the design of individual funding calls, we offer recommendations for funders and the HCI community on the design process of funding programs. The design process of grantmaking is influenced by funders’ connections to other stakeholders (Saha et al., 2022a). From our dataset, we observed that private tech companies were not only providers of AI products, resources, and training support, but also funders themselves. Through investment in AI4SG, private tech companies could not only expand their market influence (Henriksen and Richey, 2022; Lukacz, 2024; Nost and Colven, 2022) but also influence other funders’ beliefs about AI’s potential and their funding priorities. At the same time, the influence of communities on funding documents remains scant (c.f.  (Saha et al., 2022a)). To reorient AI4SG towards a more balanced approach, we see the need for funders to engage with intended communities and the HCI community (Section 5.3.1). We further outline implications for HCI researchers and practitioners to support the AI4SG funding design process in Section 5.3.2.
@@ 5.3.1. For AI4SG funders: engage with intended communities
Engaging with communities in the design of funding calls would enable funders to align their agendas with the needs of those communities early in the decision-making process, before projects even begin (Saha et al., 2022a; Balestrini et al., 2014). For funding programs that target projects in a region far from the funding organizations’ locations, it is essential to have regional partners with a more solid contextual understanding, such as regional research hubs. They can serve as “intermediaries” between funding organizations and intended end beneficiaries that funding organizations don’t have direct access to (Dell and Kumar, 2016; Saha et al., 2022a).
However, there are several challenges that funders might face in engaging with communities. Gardner et al. point out that staff in funding organizations don’t necessarily have internal expertise in ethical AI development to craft and incorporate responsible AI principles in public AI funding (Gardner et al., 2022). Relatedly, through interviews with decision-makers from grant foundations for ICTD projects, Saha et al. found that while funders often have a desire to connect with communities, they struggle to incorporate community voices into the grantmaking process, due to barriers such as geographical distances, language barriers, and a lack of institutional incentives (Saha et al., 2022a). Engaging with the HCI community could help address some of these challenges in incorporating community voices into the grantmaking process.
@@ 5.3.2. For the HCI community: implications for practice and research
While there is a rich body of resources for community-collaborative approaches (CCA) (Cooper et al., 2022), there exist gaps for the HCI community to bridge between frameworks and practices, such as implementing formal evaluations of participatory design projects (Bossen et al., 2016; Nguyen, 2022). One possible reason for these gaps is the lack of HCI expertise present in non-academic institutions. HCI practitioners, who work directly with non-academic stakeholders to implement frameworks, can establish partnerships with or apply for positions in funding organizations to put CCA frameworks into practice (such as consulting contracts or sabbatical positions at funding organizations like the U.S. NSF).
Specifically, HCI practitioners can support funders in identifying key stakeholders to engage with in communities, designing training materials on community engagement for project teams, and providing community engagement workshops for both the project teams and the funders. Moreover, the HCI practitioners can help foster reflexivity in the grantmaking process, such as reflection workshops that bring funders, regional partners, representatives of intended beneficiaries, and targeted applicants together to learn from past funding program experiences, leveraging deliberation tools when considering strategies around AI (such as the Situated AI Guidebook  (Kawakami et al., 2024b)). Wellcome Trust’s design process of Mental Health Data Prizes (C8 and C9) exemplifies the benefits of feedback workshops (Section 4.4.2).
@@
Our analyses also surface opportunities for future HCI research to further understand the role of funders in AI4SG initiatives, as well as pathways to shifting towards more balanced practices. For example, one promising direction is to interview staff members of AI4SG funding organizations to better understand the motivations, internal deliberations, and tensions between what is articulated in funding documents and the project selection process; our document analysis can serve as a basis for developing interview protocols for funders, potential applicants, and funded teams (Bowen, 2009). Future work could also engage with funded project teams to better understand their funding application strategies and needs (c.f., (Foster et al., 2024; Baxter et al., 2016)). Linking funding documents and funded project outcomes could be fruitful in identifying alignments or gaps between visions communicated through funding rhetoric and practices on the ground. Future work could also categorize commonalities of success and failure specific to AI4SG initiatives (similar to Dodson et al. (2012)’s analysis of ICTD projects). Since publicly available written materials may be biased toward success stories, focusing on one or a few case studies could yield insightful analysis. Last but not least, future work could expand the analysis to AI4SG funding documents published in languages other than English to identify funders located outside of Western societies and examine their funding agendas.
Furthermore, the HCI research community can offer counter-narratives to the more techno-centric imaginaries of AI in funding calls, such as those that emphasize diverse ways of knowing and sustainable development of AI within planetary boundaries. Some emerging critical reassessments of current AI design practices and public discourse around AI include work on AI abandonment (Johnson et al., 2024), post-capitalist AI systems based on feminist values (Varon et al., 2024), Indigenous thoughts (Ofosu-Asare, 2024; Lewis et al., 2024), post-colonial computing (Sultana and Ahmed, 2019), and post-growth philosophies (Sharma et al., 2023; Kuijer and Laschke, 2024). HCI scholars can make these perspectives more accessible to funders and AI4SG stakeholders beyond the research community by translating them into public-facing texts, such as articles in ACM Interactions magazine and online posts. These textual instruments could help bridge academic insights and real-world decision-making.
@@ 6. Conclusion
Artificial Intelligence for Social Good (AI4SG) emerges as a growing body of research and practice that explores the potential of AI and data-driven technologies to tackle social problems like public health (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). A growing body of work shows that AI4SG initiatives face various challenges in real-world contexts, including a lack of community engagement and project sustainability (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). While prior work has focused on the roles that project teams—typically AI developers and their nonprofit partners–play in project outcomes, an emerging body of work points to the significant influence of funding agendas in shaping the direction, design, and priorities of technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Gardner et al., 2022).

### v3

@@ 5.2.1. Background: situating AI4SG in well-scoped contexts
Funding calls should orient towards projects that address well-scoped and contextually relevant problems. This doesn’t mean that funders should pre-define the problems, but they can support applicants by offering examples of relevant AI use cases or existing literature review on the social domain (Section 4.1.2). It also involves requiring the active participation of communities during the problem formulation and solution brainstorming stage, where applicants and communities collaboratively investigate which technologies are appropriate for their specific problem. In a systematic review of community collaborative approaches, Cooper et al. (2022) highlight that many projects involve communities only during limited stages (e.g., delivery of activities), rather than throughout the full research process—including problem framing and data analysis. Yet, as Bratteteig and Wagner (2016) argue, engaging communities in deciding which problems to address is key, not only for inclusion but for redistributing power and ensuring they have real influence over project direction and outcomes. This level of participation reflects a deeper form of agency than simply evaluating or implementing pre-defined solutions (Delgado et al., 2023).
@@ 5.2.2. Expected Outcomes: defining success with intended communities
Successful outcomes in AI4SG projects should be defined collaboratively with the communities they aim to serve. Community perspectives on success might differ from funders’ or researchers’ goals, which tend to emphasize metrics like model accuracy or academic output (Lin et al., 2024; Erete et al., 2023). For example, Lin et al. (2024) found that community organizations in AI4SG partnerships value technical capacity building and project visibility more than academic output. Aligning success metrics early on helps manage expectations and reduces the risk of disappointment as projects progress (Taylor et al., 2013). While stakeholder goals may still sometimes conflict (Ho et al., 2009a), projects could aim for pre-determined mutual benefit—for instance, through fair compensation for communities’ time and expertise or by supporting informal learning opportunities (Carroll and Rosson, 2013). This would also require attention to non-traditional (and often not quantifiable) metrics that capture relationship quality (Cooper et al., 2022; Erete et al., 2023). For example, funders could request letters of support and experiences from communities as part of the project’s evaluation, alongside the quantitative evaluation of AI systems.
When real-world deployment is the goal, it is essential to provide funding support and maintenance resources post-deployment, which is notably absent from most funding programs in our dataset. In categorizing ICTD projects’ outcomes, Heeks (2002) calls attention to the high rate of projects that might have initially succeeded but break down over time. As Jackson (2014) argues, “repair and maintenance” is a critical part of technology design, yet it pervasively falls out of view when the dominant discourse of technology focuses on innovations as a production, and not a process. The maintenance work of a technology system is not only necessary for sustaining the system that will inevitably keep breaking (Jackson, 2014), it also builds relationships among workers who collaborate on the practical work (Irani and Silberman, 2014; D’Ignazio, 2024). Sustainable maintenance can involve building local capacity and intentionally transferring technology ownership to communities. For example, the research team behind SignSupport, a mobile app co-developed with the Deaf Community of Cape Town to support pharmacy interactions, trained community members to manage a computer lab and take on system administration roles, while researchers stepped back to allow community leadership to grow without interference (Dearden and Tucker, 2016).
@@
Funding calls should be structured to include intended community members as co-leaders or co-applicants, ensuring they have the power to shape project direction from the outset—and importantly, that they receive funding directly. As Bratteteig and Wagner (2016) point out, the “skill composition” of a design team significantly shapes how problems and solutions are defined. Including community members in core project roles enables them to define the problem space and success criteria, and ensures their labor and expertise are properly acknowledged and compensated (Pine et al., 2020). This approach also helps redistribute power within AI4SG collaborations, supporting more equitable relationships between communities and research institutions (Erete et al., 2023). Crucially, when engaging communities throughout the project, it is important to recognize and compensate for the various forms of labor they contribute—from sharing lived experience and domain expertise to facilitating connections within their communities. Future HCI research could work towards quantifying both visible and invisible labor (modeling after recent work in, for example, crowd work (Toxtli et al., 2021) and online volunteer moderation (Li et al., 2022)), while also acknowledging that some forms of labor in community engagement, such as emotional labor, may be difficult or inappropriate to quantify.
@@ 5.2.4. Funders’ Support: valuing community engagement as a necessary part of AI4SG work
Funders can better support community engagement in AI4SG by recognizing that relationship-building is a foundational phase of the work, which often occurs before a clear project direction is even established (Carroll and Rosson, 2013; Cooper et al., 2022). This phase is essential to community-collaborative approaches (Cooper et al., 2022), and it requires the regular presence of project teams in meetings or other community contexts, which takes time and emotional labor (Dearden and Tucker, 2016). Long-term relationships enable shared agendas to emerge, with community priorities guiding technology development from the outset (Dearden and Tucker, 2016; Carroll and Rosson, 2013). One concrete step funders can take is to extend the typical AI4SG funding timeline (in our corpus: with a median of one and a half years) to accommodate this early-stage work. Relationship-building alone can take months (Erete et al., 2023) or even years (Carroll and Rosson, 2013; Dearden and Tucker, 2016), and should be supported as a formal project phase, even prior to defining problems or solutions. One key challenge is determining appropriate timelines and support for this work. Future research in community collaborative approaches could analyze researchers’ reflections (e.g., (Erete et al., 2023; Dearden and Tucker, 2016; Carroll and Rosson, 2013)) to better understand the duration and nature of relationship-building across projects.
@@
Furthermore, funders should also treat community engagement as a critical skill that requires training and support, on par with technical expertise. Many AI practitioners are trained in Computer Science departments that don’t routinely provide training in community engagement. As Okolo and Lin (2024) observe, AI4SG practitioners often use a hodgepodge of ad-hoc approaches to engage with end users. Funders should recognize this knowledge gap and provide clear guidance to project teams on how to properly identify stakeholders and engage with end beneficiaries. There are multiple places where this guidance could be provided: as part of the funding call itself or additional training support that many funders already provide. Funders could consider citing existing guides in community engagement from the HCI community, such as the Equity-Centered Community Design Field Guide (Creative Reaction Lab, 2024), the Atlanta Community Engagement Playbook (Taranji Alvarado et al., 2024), and “readiness for partnership” checks (Pine et al., 2020). Other design toolkits, like the Envisioning Cards (Friedman et al., 2024), might also be useful for AI4SG project teams to identify and engage with stakeholders. Since AI introduces additional hurdles to community-based designs, including requiring intensive data work (Miceli and Posada, 2022; Passi and Sengers, 2020; Sambasivan et al., 2021) and transparency issues (Okolo and Lin, 2024), HCI researchers could further support funders by tailoring existing toolkits to the specific contexts of AI4SG, building on emerging Participatory AI4SG frameworks like the PACT framework (Bondi et al., 2021).
@@ 5.3. Calling for Meaningful Engagement: Recommendations for the Funding Design Process
Beyond the design of individual funding calls, we offer recommendations for funders and the HCI community on the design process of funding programs. The design process of grantmaking is influenced by funders’ connections to other stakeholders (Saha et al., 2022a). From our dataset, we observed that private tech companies were not only providers of AI products, resources, and training support, but also funders themselves. Through investment in AI4SG, private tech companies not only expand their market influence (Henriksen and Richey, 2022; Lukacz, 2024; Nost and Colven, 2022) but also influence other funders’ beliefs about AI’s potential and their funding priorities. At the same time, the influence of communities on funding documents remains scant (c.f. (Saha et al., 2022a)). To reorient AI4SG towards a more balanced approach, we see the need for funders to engage with intended communities and the HCI community (Section 5.3.1). We further outline implications for HCI researchers and practitioners to support the AI4SG funding design process in Section 5.3.2.
@@ 5.3.1. For AI4SG funders: engage with intended communities
Engaging with communities in the design of funding calls would enable funders to align their agendas with the needs of those communities early in the decision-making process, before funding programs are even open for applications (Saha et al., 2022a; Balestrini et al., 2014). For funding programs that target projects in a region far from the funding organizations’ locations, it is essential to have regional partners with a more solid contextual understanding, such as regional research hubs and local community organizations. They can serve as “intermediaries” between funding organizations and intended end beneficiaries that funding organizations don’t have direct access to (Dell and Kumar, 2016; Saha et al., 2022a).
However, several challenges may arise when funders engage with communities. Gardner et al. (2022) point out that staff in funding organizations may not have internal expertise in ethical AI development to craft and incorporate responsible AI principles into public AI funding. Relatedly, through interviews with decision-makers from grant foundations for ICTD projects, Saha et al. (2022a) found that while funders often have a desire to connect with communities, they struggle to incorporate community voices into the grantmaking process, due to barriers such as geographical distances, language barriers, and a lack of institutional incentives. Engaging with the HCI community could help address some of these challenges in incorporating community voices into the grantmaking process.
@@ 5.3.2. For the HCI community: implications for practice and research
While there is a rich body of resources for community collaborative approaches (Cooper et al., 2022), there exist gaps for the HCI community to bridge between frameworks and practices, such as implementing formal evaluations of participatory design projects (Bossen et al., 2016; Nguyen, 2022). One possible reason for these gaps is the lack of HCI expertise present in non-academic institutions. HCI practitioners, who work directly with non-academic stakeholders to implement frameworks, can establish partnerships with or apply for positions at funding organizations to put community collaborative approaches into practice (such as consulting contracts or sabbatical positions at funding organizations like the National Science Foundation in the U.S.).
Specifically, HCI practitioners can support funders in identifying key stakeholders to engage with in communities, designing training materials on community engagement for project teams, and providing community engagement workshops for both the project teams and the funders. Moreover, HCI practitioners can help foster reflexivity in the grantmaking process, such as reflection workshops that bring funders, regional partners, representatives of intended beneficiaries, and targeted applicants together to learn from past funding program experiences, leveraging deliberation tools when considering strategies around AI (such as the Situated AI Guidebook (Kawakami et al., 2024b)). Wellcome Trust’s design process of Mental Health Data Prizes (C8 and C9) exemplifies the benefits of feedback workshops (Section 4.4.2).
@@
Our analyses also surface opportunities for future HCI research to further understand the role of funders in AI4SG initiatives, as well as pathways to shifting towards more balanced practices. We list a few promising research directions below:
  • One promising direction is to interview staff members of AI4SG funding organizations to better understand the motivations, internal deliberations, and tensions between what is articulated in funding documents and the project selection process. Our document analysis can serve as a basis for developing interview protocols for funders, potential applicants, and funded teams (Bowen, 2009).
  • Future work could also engage with funded project teams to better understand their funding application strategies and needs (c.f., (Foster et al., 2024; Baxter et al., 2016)).
  • Linking funding documents to funded project outcomes could be fruitful for identifying alignments or gaps between the visions communicated through funding rhetoric and practices on the ground.
  • Future work could also categorize commonalities of success and failure specific to AI4SG initiatives (similar to Dodson et al. (2012)’s analysis of ICTD projects). Since publicly available written materials may be biased toward success stories, focusing on one or a few case studies could yield insightful analysis.
  • Future work could expand the analysis to include AI4SG funding documents published in languages other than English to identify funders outside Western societies and examine their funding agendas.
Furthermore, the HCI research community can offer counter-narratives to the more techno-centric imaginaries of AI in funding calls, such as those that emphasize diverse ways of knowing and sustainable development of AI within planetary boundaries. Some emerging critical reassessments of current AI design practices and public discourse around AI include work on AI abandonment (Johnson et al., 2024), post-capitalist AI systems based on feminist values (Varon et al., 2024), Indigenous thoughts (Ofosu-Asare, 2024; Lewis et al., 2024), post-colonial computing (Sultana and Ahmed, 2019), and post-growth philosophies (Sharma et al., 2023; Kuijer and Laschke, 2024). HCI scholars can make these perspectives more accessible to funders and AI4SG stakeholders by translating them into public-facing texts, such as articles in ACM Interactions magazine and online posts. These textual instruments could help bridge the gap between academic insights and real-world decision-making.
@@ 6. Conclusion
Artificial Intelligence for Social Good (AI4SG) emerges as a growing body of research and practice that explores the potential of AI and data-driven technologies to tackle social problems like public health (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). A growing body of work shows that AI4SG initiatives face various challenges in real-world contexts, including deployment and project sustainability (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). While prior work has focused on the roles that project teams—typically AI developers and their nonprofit partners–play in project outcomes, an emerging body of work points to the significant influence of funding agendas in shaping the direction, design, and priorities of technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Gardner et al., 2022).
@@
Acknowledgements.
We are grateful to the anonymous reviewers, Weiwei Pan, Eura Shin, and members of the Intelligent Interactive Systems Group at Harvard and the Data + Feminism Lab at MIT for their valuable feedback on the early drafts of this work.